OpenStudy (anonymous):

Solve each equation. Write a justification for each step. 16. 8 - 5s = 1 17. 0.4t + 3 = 1.6 18. 38 = -3w + 2

OpenStudy (anonymous):

-5s = 1 - 8 s = 1 - 8 / -5

OpenStudy (anonymous):

\[16.8-5s=1\rightarrow16.8-1=5s \rightarrow 5s=15.8\rightarrow 5s=15+\frac{4 }{ 5 }\]\[s=\frac{ 15 }{ 5 }+\frac{ \frac{ 4 }{ 5 } }{ 5 }=3+\frac{ 4 }{ 25 }=\frac{ 79 }{ 25 }=3.16\] And, wait, never mind, I just realized the 16. 17. and 18. are the numbers of the questions...

OpenStudy (anonymous):

\[8-5s = 1 \rightarrow1-8=-5s \rightarrow8-1=5s \rightarrow 5s=7\rightarrow s=\frac{ 7 }{ 5 }\]\[0.4t+3=1.6\rightarrow0.4t=1.6-3\rightarrow0.4t=-1.4\rightarrow t=\frac{ -1.4 }{ 0.4 }\rightarrow t=-\frac{ 7 }{ 2 }\] \[38=3w+2\rightarrow38-2=3w \rightarrow36=3w \rightarrow w=\frac{ 36 }{ 3 }\rightarrow w=12\]